# Visual Autoregressive Modeling:  Scalable Image Generation via Next-Scale Prediction

### v1

@@ 3.3 Implementation details
4 Empirical Results
This section first compares VAR with other image generative model families [^3] For fairness, methods with advanced VQVAE are excluded in this model family comparison, e.g. the BERT-style masked-pred model MagViT-2 [74]. Practitioners can combine them with VAR for better results. in terms of performance and and efficiency in Sec. 4.1. Evaluations on the scalability and generalizability of VAR models are presented in Sec. 4.2 and Sec. 4.3. We do some ablations and visualizations at the end.
@@ 4.1 State-of-the-art image generation
Compared with popular diffusion transformer.444[46] do not report DiT-L/2’s performance with CFG so we use official code to reproduce it.  The VAR model surpasses the recently popular diffusion models Diffusion Transformer (DiT), which serves as the precursor to the latest Stable-Diffusion 3 [18] and SORA [8], in multiple dimensions: 1) In image generation diversity and quality (FID and IS), VAR with 2B parameters consistently performs better than DiT-XL/2 [46], L-DiT-3B, and L-DiT-7B [2]. VAR also maintains comparable precision and recall. 2) For inference speed, the DiT-XL/2 requires 45$\times$ the wall-clock time compared to VAR, while 3B and 7B models [2] would cost much more. 3) VAR is considered more data-efficient, as it requires only 350 training epochs compared to DiT-XL/2’s 1400. 4) For scalability, Fig. 3 and Tab. 1 show that DiT only obtains marginal or even negative gains beyond 675M parameters. In contrast, the FID and IS of VAR are consistently improved, aligning with the scaling law study in Sec. 4.2. These results establish VAR as a more efficient and scalable model for image generation than models like DiT.
@@ 4.2 Power-law scaling laws
Scaling laws with model parameters $N$.  We first investigate the test loss trend as the VAR model size increases. The number of parameters $N(d)=73728\,d^{3}$ for a VAR transformer with depth $d$ is specified in (8). We varied $d$ from $6$ to $30$, yielding 12 models with 18.5M to 2.0B parameters. We assessed the final test cross-entropy loss $L$ and token prediction error rates $Err$ on the ImageNet validation set of 50,000 images [14]. We computed $L$ and $Err$ for both the last scale (at the last next-scale autoregressive step), as well as the global average. The results are plotted in Fig. 5, where we observed a clear power-law scaling trend for $L$ as a function of $N$, as consistent with [30, 22, 27, 1]. The power-law scaling laws can be expressed as:
@@
$L_{\text{last}}=(2.2\cdot 10^{-5}C_{\text{min}})^{-0.13}$ $\text{and}~{}L_{\text{avg}}=(1.5\cdot 10^{-5}C_{\text{min}})^{-0.16},$ (13)
$Err_{\text{last}}=(8.1\cdot 10^{-2}C_{\text{min}})^{-0.0067}$ $\text{and}~{}Err_{\text{avg}}=(4.4\cdot 10^{-2}C_{\text{min}})^{-0.011}.$ (14)
These relations (13, 14) hold across 6 orders of magnitude in $C_{\text{min}}$, and our findings are consistent with those in [30, 22]: when trained with sufficient data, larger VAR transformers are more compute-efficient because they can reach the same level of performance with less computation.
Visualizations.  To better understand how VAR models are learning when scaled up, we compare some generated $256\times 256$ samples from VAR models of 4 different sizes (depth 6, 16, 26, 30) and 3 different training stages (20%, 60%, 100% of total training tokens) in Fig. 7. To keep the content consistent, a same random seed and teacher-forced initial tokens are used. The observed improvements in visual fidelity and soundness are consistent with the scaling laws, as larger transformers are thought able to learn more complex and fine-grained image distributions.
@@
4.3 Zero-shot task generalization
@@
4.4 Ablation Study
Component-wise ablation.  We further test some key components in VAR. By replacing the standard Layer Normalization (LN) with Adaptive Layer Normalization (AdaLN), VAR starts yielding better FID than baseline. By using the top-$k$ sampling similar to the baseline, VAR’s FID is further improved. By using the classifier-free guidance (CFG) with ratio $2.0$, we reach the FID of 3.60, which is 15.05 lower to the baseline, and its inference is still 45 times faster. Due to the observed effectiveness, we equip our final VAR models with AdaLN, top-$k$ sampling, and classifier-free guidance. We finally scale up VAR size to 2.0B and achieve an FID of 1.80. This is 16.85 better than the baseline FID.
Table 3: Ablation study of VAR. The first two rows compare GPT-2-style transformers trained with AR or VAR algorithm. Subsequent lines show the influence of VAR enhancements. “AdaLN”: adaptive layernorm. “CFG”: classifier-free guidance. “Cost”: inference cost relative to the baseline. “$\Delta$”: reduction in FID.
@@ 4.4 Ablation Study
6 $+$Scale up 2.0B VAR-$d$30 600 2.0 0.052 1.80 --16.85
5 Future Work
@@ Appendix B Time complexity of AR and VAR generation
Figure 10: Model comparison on ImageNet 256$\times$256 benchmark.
Figure 11: More ImageNet 256$\times$256 generation samples.

### v2

@@ 5.1 State-of-the-art image generation
Compared with popular diffusion transformer.  The VAR model surpasses the recently popular diffusion models Diffusion Transformer (DiT), which serves as the precursor to the latest Stable-Diffusion 3 [29] and SORA [14], in multiple dimensions: 1) In image generation diversity and quality (FID and IS), VAR with 2B parameters consistently performs better than DiT-XL/2 [63], L-DiT-3B, and L-DiT-7B [3]. VAR also maintains comparable precision and recall. 2) For inference speed, the DiT-XL/2 requires 45$\times$ the wall-clock time compared to VAR, while 3B and 7B models [3] would cost much more. 3) VAR is considered more data-efficient, as it requires only 350 training epochs compared to DiT-XL/2’s 1400. 4) For scalability, Fig. 3 and Tab. 1 show that DiT only obtains marginal or even negative gains beyond 675M parameters. In contrast, the FID and IS of VAR are consistently improved, aligning with the scaling law study in Sec. 5.2. These results establish VAR as potentially a more efficient and scalable model for image generation than models like DiT.
@@ 5.2 Power-law scaling laws
An appealing phenomenon is that both [43] and [35] never observed deviation from these linear relationships at the higher end of $X$, although flattening is inevitable as the loss approaches zero.
@@
Scaling laws with model parameters $N$.  We first investigate the test loss trend as the VAR model size increases. The number of parameters $N(d)=73728\,d^{3}$ for a VAR transformer with depth $d$ is specified in (8). We varied $d$ from $6$ to $30$, yielding 12 models with 18.5M to 2.0B parameters. We assessed the final test cross-entropy loss $L$ and token prediction error rates $Err$ on the ImageNet validation set of 50,000 images [24]. We computed $L$ and $Err$ for both the last scale (at the last next-scale autoregressive step), as well as the global average. Results are plotted in Fig. 5, where we
observed a clear power-law scaling trend for $L$ as a function of $N$, as consistent with [43, 35, 38, 1]. The power-law scaling laws can be expressed as:
@@
$L_{\text{last}}=(2.2\cdot 10^{-5}C_{\text{min}})^{-0.13}$ (13)
$L_{\text{avg}}=(1.5\cdot 10^{-5}C_{\text{min}})^{-0.16},$ (14)
$Err_{\text{last}}=(8.1\cdot 10^{-2}C_{\text{min}})^{-0.0067}$ (15)
$Err_{\text{avg}}=(4.4\cdot 10^{-2}C_{\text{min}})^{-0.011}.$ (16)
These relations (14, 16) hold across 6 orders of magnitude in $C_{\text{min}}$, and our findings are consistent with those in [43, 35]: when trained with sufficient data, larger VAR transformers are more compute-efficient because they can reach the same level of performance with less computation.
5.3 Visualization of scaling effect
To better understand how VAR models are learning when scaled up, we compare some generated $256\times 256$ samples from VAR models of 4 different sizes (depth 6, 16, 26, 30) and 3 different training stages (20%, 60%, 100% of total training tokens) in Fig. 7. To keep the content consistent, a same random seed and teacher-forced initial tokens are used. The observed improvements in visual fidelity and soundness are consistent with the scaling laws, as larger transformers are thought able to learn more complex and fine-grained image distributions.
@@ 5.3 Visualization of scaling effect
6 Zero-shot task generalization
@@
Table 3: Ablation study of VAR. The first two rows compare GPT-2-style transformers trained under AR or VAR algorithm without any bells and whistles. Subsequent lines show the influence of VAR enhancements. “AdaLN”: adaptive layernorm. “CFG”: classifier-free guidance. “Attn. Norm.”: normalizing $q$ and $k$ to unit vectors before attention. “Cost”: inference cost relative to the baseline. “$\Delta$”: FID reduction to the baseline.
@@
5 $+$Attn. Norm. 310M VAR-$d$16 600 2.0 0.022 3.30 --15.35
6 $+$Scale up 2.0B VAR-$d$30 600 2.0 0.052 1.73 --16.85
7 Ablation Study
Component-wise ablation.  We further test some key components in VAR. By replacing the standard Layer Normalization (LN) with Adaptive Layer Normalization (AdaLN), VAR starts yielding better FID than baseline. By using the top-$k$ sampling similar to the baseline, VAR’s FID is further improved. By using the classifier-free guidance (CFG) with ratio $2.0$ and normalizing $q$ and $k$ to unit vectors before attention, we reach the FID of 3.30, which is 15.35 lower to the baseline, and its inference speed is still 45 times faster. We finally scale up VAR size to 2.0B and achieve an FID of 1.73. This is 16.85 better than the baseline FID.
8 Limitations and Future Work
@@ Appendix B Time complexity of AR and VAR generation
Figure 10: Model comparison on ImageNet 256$\times$256 benchmark. More generated 512$\times$512 samples by VAR can be found in the submitted Supplementary Material zip file.
Figure 11: Some generated 256$\times$256 samples by VAR trained on ImageNet. More generated 512$\times$512 samples by VAR can be found in the submitted Supplementary Material zip file.